OpenStudy (anonymous):

Let f(x) = 9x – 2 and g(x) = –x + 3. Find f(g(x)).

OpenStudy (anonymous):

f(g(x))=f(-x+3)=9(-x+3)-2=-9x+27-2= -9x+25